# Electrodynamics of Vortices in Quasi-2D Scalar Bose-Einstein Condensates

### v3

@@ II.2 Case 2: beyond the Gross-Pitaevskii equation
where $G\left(\bm{r},t\right)$ and $\bm{F}\left(\bm{r},t\right)$ are real functions to be determined by the equation $i\hbar\partial\psi\left(\bm{r},t\right)/\partial t$ in the model under consideration. Said differently, they are model-dependent. We shall focus on the specific forms of $\bm{F}\left(\bm{r},t\right)$ or $G\left(\bm{r},t\right)$ for a rotating quasi-2D scalar BEC without dissipation in section II.2.1. Nevertheless, in section IV, we will show that one can still build the effective Maxwell’s equations without specifying them.
@@
It can be shown that so far, every result in this section II is also valid in 3D space. In the following subsection, we will consider a rotating quasi-2D scalar BEC without dissipation as an example.
@@ II.2.1 Example: rotating quasi-2D scalar BEC without dissipation
Motivated by Ref. [68], we will consider a rotating quasi-2D scalar BEC with angular velocity $\bm{\Omega}_{\perp}\left(\bm{r},t\right)=\Omega_{\perp}\left(\bm{r},t\right)\bm{e}_{\perp}$ that depends on position or time. In the rotating frame, the GPE in Eq. (6) becomes [55]
@@ III Vortex quantization
where $\delta\left(\bm{r}\right)$ is the Dirac delta function, and $\bm{r}=\bm{r}_{\alpha_{j}}\left(t\right)$ is the position of the core of the vortex with topological charge $q_{j}$ at time $t$. In BEC experiments or numerical simulations, the typical size of the vortex is about the order of the healing length (coherence length) $\xi_{h}\coloneqq\hbar/\sqrt{2Mgn_{m}}$, where $n_{m}$ is the mean BEC number density [39].
In general, even for non-rotating BEC, the topological charge of the vortex may change over time [69]. It may flip sign (e.g., from $q=1$ to $q=-1$), turning a vortex into an antivortex, in a non-rotating quasi-2D scalar BEC under anisotropic trap potential and small $g$ [70, 69]. The non-conservation of the number of vortices is observed in non-rotating scalar bosons under the second-order phase transition [12, 13, 14, 15]. Its growth can be explained via the Kibble-Zurek mechanism [5], while its decay can result from coarsening [71, 72]. Also, this change in the number of vortices is numerically shown in the stochastic GPE [11] and in the stochastic projected GPE [73].
Considering the case of time-dependent topological charges, for multiple vortices in the non-rotating quasi-2D scalar BEC in the region $\mathcal{A}$, Eqs. (29) can be generalized to
@@
and for special cases where $n\left(\bm{r},t\right)=0$ only at cores of vortices,
@@
implying that the quantum curl dynamics can describe the system. Though it is not always possible to build effective Maxwell’s equations in the quantum curl dynamics [74], Eq. (33) will be used in section IV to define the effective free electric current density in the effective Maxwell’s equations, and in section VI to present how it is related to the time change of the circulation of $\bm{v}_{s}\left(\bm{r},t\right)$.
@@ IV The connection between vortices in quasi-2D BEC and electrodynamics
The mathematical connection between static vortices in a non-rotating quasi-2D scalar BEC and 2D electrostatics within the GPE and PVM is well established [29]. In this section, we will derive the duality between vortices in a quasi-2D BEC and 2D electrodynamics in general. Motivated by Ref. [37], let us define the vortex charge density $\rho_{v}\left(\bm{r},t\right)$ such that
@@
We define $\rho_{v}\left(\bm{r},t\right)$ in this way because the curl of $\bm{v}_{s}\left(\bm{r},t\right)$ doesn’t need to be the Dirac delta function if the superfluid density $n\left(\bm{r},t\right)$ is zero within some finite region around the core of the vortex. For example, the nonlinearity of the field equation for superfluid ${}^{4}\textrm{He}$ changes the curl of the superfluid velocity $\bm{v}_{s}\left(\bm{r},t\right)$ to the smeared delta function (see Eq. (1.64) in Part II of Ref. [32]). Note that setting $\rho_{v}\left(\bm{r},t\right)=\sum_{j=1}^{\infty}q_{j}\left(t\right)\delta\left(\bm{r}-\bm{r}_{\alpha_{j}}\left(t\right)\right)$ corresponds to the PVM, so we also consider beyond the PVM.
@@
However, Eq. (42) does not imply vortex charge conservation; the vortex charge may not be conserved as $q\left(t\right)$ may depend on time $t$. This consideration, including the non-conservation of the vortex charge, is one of the main differences between our work and the previous works relying on the conservation of the vortex charge [30, 67, 75, 76]. In section VI, Eq. (42) will be used to show how to calculate the change of the circulation of $\bm{v}_{s}\left(\bm{r},t\right)$ in time.
@@
If $U_{\rm sf}\left(\bm{r},t\right)$ is constant in space, Eq. (43) cannot be distinguished from Maxwell’s equations with the effective magnetic field strength $\bm{H}_{\rm sf}\left(\bm{r},t\right)=0$. Therefore, let us define $\bm{H}_{\rm sf}\left(\bm{r},t\right)$ as
@@
From the physical point of view, we will set $c_{\rm sf}$ to be the maximum value of the speed of sound in the scalar BEC, i.e., $c_{s}\coloneqq\sqrt{gn_{\rm max}/M}$, where $n_{\rm max}$ is the maximum value of $n\left(\bm{r},t\right)$ [44].
@@
where the effective free magnetic charge density $\rho_{m,\rm{sf}}\left(\bm{r},t\right)$ is zero (see Eq. (47)). Equation (51) implies the conservation of the effective free magnetic charge, so the effective free magnetic charge always remains zero.
@@
implying that the vortex (free electric charge) moves parallel to $\bm{v}_{P}\left(\bm{r},t\right)$, not parallel to the superfluid velocity $\bm{v}_{s}\left(\bm{r},t\right)$ in general as is suggested in Ref. [54] when there is dissipation. The obtained duality is summarized in Fig. 1 and Table 3.
Figure 1: Duality between the vortices in the non-rotating quasi-2D scalar BEC and the electrodynamics in the matter. The effective Maxwell’s equations are shown in Table 3.
Table 3: Effective Maxwell’s equations for the vortices in the non-rotating quasi-2D scalar BEC. $\rho_{v}\left(\bm{r},t\right)$ is the vortex charge density defined in Eq. (35), $\bm{J}_{\rm sf}\left(\bm{r},t\right)$ is the effective free electric current density defined in Eq. (41), $\bm{J}_{m,\rm{sf}}\left(\bm{r},t\right)$ is the effective free magnetic current density defined in Eq. (49), and $c_{\rm sf}$ is the maximum value of the sound of speed in the scalar BEC. Refer to Eqs. (36), (38), (44) for the definitions of the effective fields.
@@
We emphasize that our results are of broader generality than those reported in Refs. [36, 37], that assumed a uniform condensate density with negligible fluctuation in a non-rotating system, or Ref. [38], that assumed inhomogeneous time-independent condensate density in a non-rotating system to derive Maxwell’s equations for the (2+1) dimensional superfluid universe. By contrast to these preceding works, the duality described in this section IV can also be applied in the case of an inhomogeneous time-dependent condensate density $n\left(\bm{r},t\right)$ in the (2+1) dimensional spacetime, even in the rotating frame (refer to Eqs. (26) and (28) for a rotating quasi-2D scalar BEC without dissipation). This extension to the inhomogeneous time-dependent $n\left(\bm{r},t\right)$ is important to study dynamics of vortices because (i) vortices can move, and (ii) $n\left(\bm{r},t\right)=0$ at the core of the vortex. Due to those properties, fluctuation of $n\left(\bm{r},t\right)$ cannot be neglected in general.
@@
The results in this section show that one may consider vortices in a quasi-2D scalar BEC as free electric charges in 2D matter. However, there is one thing missing in the duality we presented: what is the effective magnetic field $\bm{B}_{\rm sf}\left(\bm{r},t\right)$? In the next section, we will define the direction of $\bm{B}_{\rm sf}\left(\bm{r},t\right)$ from the effective Lorentz force and show that the phenomenological damped PVM [54] can be derived.
@@ V The effective force acting on vortices
From the duality we found above, one can see that the effective Lorentz force per unit area acting on a vortex, that is, the “effective free electric charge”, has two contributions $\bm{f}_{v}\left(\bm{r},t\right)=\bm{f}_{1}\left(\bm{r},t\right)+\bm{f}_{2}\left(\bm{r},t\right)$ where
@@
Note that Eq. (57) is consistent with the “force per unit length on a vortex line” in Refs. [64, 30]. Another notable thing is that Eq. (60) is similar to the phenomenological damped-PVM [54]. For a non-rotating quasi-2D scalar BEC confined in a box-like trapping potential ($V\left(\bm{r},t\right)=0$) as considered in Ref. [54], $M\partial\bm{v}_{s}\left(\bm{r},t\right)/\partial t=\bm{f}_{K}\left(\bm{r},t\right)+\bm{F}_{\rm sf}\left(\bm{r},t\right)-g\nabla n\left(\bm{r},t\right)$, where $\bm{F}_{\rm sf}\left(\bm{r},t\right)$ is the additional force due to the dissipation (refer to section II.2). In the GPE for a non-rotating quasi-2D BEC, minimization of the mean-field energy with respect to the phase of the mean-field $\psi\left(\bm{r},t\right)$ gives $\nabla\cdot\bm{v}_{s}\left(\bm{r},t\right)=0$ [32]. As $\nabla\cdot\bm{P}_{\rm sf}\left(\bm{r},t\right)=0$ in the PVM (see Eq. (40)), for small dissipation where $\nabla\cdot\bm{v}_{s}\left(\bm{r},t\right)\simeq 0$ is still a good approximation, one may set $\bm{P}_{\rm sf}\left(\bm{r},t\right)\simeq c_{1}\left(t\right)\bm{v}_{s}\left(\bm{r},t\right)$, with $c_{1}\left(t\right)$ being some function that only depends on time $t$. Therefore, with a suitable choice of $M_{\rm sf}\left(\bm{r},t\right)$ and an appropriate definition of the effective vortex mass, one may derive the damped-PVM in Ref. [54]. Or one may use $\bm{P}_{\rm sf}\left(\bm{r},t\right)\simeq c_{1}\left(t\right)\bm{v}_{s}\left(\bm{r},t\right)$ together with the definition of $\bm{P}_{\rm sf}\left(\bm{r},t\right)$ in the second line in Eqs. (38) to get
@@ VI Temporal change of the circulation in a static area
As an example, let us consider the case where only one vortex is positioned at the center of the circular quasi-2D scalar BEC with circular symmetry. We thus assume no external drag or other effect that breaks circular symmetry. Due to this symmetry, $\bm{f}_{K}\left(\bm{r},t\right)$ is perpendicular to the boundary of any disk $\partial\mathcal{D}^{2}_{R}\left(0\right)$ with $R>0$, where $\mathcal{D}^{d}_{R}\left(\bm{r}_{c}\right)$ denotes a $d$-dimensional disk with radius $R$ centered at $\bm{r}_{c}$. Within the GPE, $\bm{F}_{\rm sf}\left(\bm{r},t\right)=0$ for a non-rotating system, while $\bm{F}_{\rm sf}\left(\bm{r},t\right)=M\nabla\left\{\bm{v}_{s}\left(\bm{r},t\right)\cdot\left[\bm{\Omega}_{\perp}\left(\bm{r},t\right)\times\bm{r}\right]\right\}$ in the case of a rotating system (see Eq. (28)), which will also be perpendicular to $\partial\mathcal{D}^{2}_{R}\left(0\right)$ due to the symmetry. Therefore, the vortex will be dynamically stable within the GPE description both in the rotating and non-rotating cases. Conversely, for a rotating or non-rotating quasi-2D scalar circular dissipationless BEC, if the system has circular symmetry, a single vortex cannot emerge at the center if there are no vortices initially. This is consistent with the findings in Ref. [24], where vortex creation is described at the border of the trap.

### v4

@@ II.2 Case 2: beyond the nonrotating Gross-Pitaevskii equation
where $G\left(\bm{r},t\right)$ and $\bm{F}\left(\bm{r},t\right)$ are real functions to be determined by the equation $i\hbar\partial\psi\left(\bm{r},t\right)/\partial t$ in the model under consideration. Said differently, they are model-dependent. We shall focus on the specific forms of $\bm{F}\left(\bm{r},t\right)$ or $G\left(\bm{r},t\right)$ for a rotating quasi-2D scalar BEC without dissipation in Sec. II.2.1. Nevertheless, in Sec. IV, we will show that one can still build the effective Maxwell’s equations without specifying them.
@@
It can be shown that so far, every result in this Sec. II is also valid in 3D space. In the following subsection, we will consider a rotating quasi-2D scalar BEC without dissipation as an example.
@@ II.2.1 Example: rotating quasi-2D scalar BEC without dissipation
Motivated by the experiment with two concentric counter-rotating superfluids [70], we will consider a rotating quasi-2D scalar BEC with angular velocity $\bm{\Omega}_{\perp}\left(\bm{r},t\right)=\Omega_{\perp}\left(\bm{r},t\right)\bm{e}_{\perp}$ that depends on position or time. In the rotating frame, the GPE in Eq. (6) becomes [57]
@@ III Vortex quantization
where $\delta\left(\bm{r}\right)$ is the Dirac delta function, and $\bm{r}=\bm{r}_{\alpha_{j}}\left(t\right)$ is the position of the core of the vortex with topological charge $q_{j}$ at time $t$. In BEC experiments or numerical simulations, the typical size of the vortex core is about the order of the healing length (coherence length) $\xi_{h}\coloneqq\hbar/\sqrt{2Mgn_{m}}$, where $n_{m}$ is the mean BEC number density [41].
In general, even for nonrotating BEC, the topological charge of the vortex may change over time. It may flip sign (e.g., from $q=1$ to $q=-1$), turning a vortex into an antivortex, in a nonrotating quasi-2D scalar BEC under anisotropic trap potential and small $g$ [71]. The nonconservation of the number of vortices is observed in nonrotating scalar bosons under the second-order phase transition [14, 15, 16, 17]. Its growth can be explained via the Kibble-Zurek mechanism [5], while its decay can result from coarsening [72, 73]. Also, this change in the number of vortices is numerically shown in the stochastic GPE under the periodic boundary conditions [13] and in the stochastic projected GPE [74].
Considering the case of time-dependent topological charges, for multiple vortices in the nonrotating quasi-2D scalar BEC in the region $\mathcal{A}$, Eqs. (29) can be generalized to
@@
and for special cases where $n\left(\bm{r},t\right)=0$ only at vortex cores,
@@
implying that the quantum curl dynamics can describe the system. Though it is not always possible to build effective Maxwell’s equations in the quantum curl dynamics [75], Eq. (33) will be used in Sec. IV to define the effective free electric current density in the effective Maxwell’s equations, and in Sec. VI to present how it is related to the time change of the circulation of $\bm{v}_{s}\left(\bm{r},t\right)$.
@@ IV The connection between vortices in quasi-2D BEC and electrodynamics
The mathematical connection between static vortices in a nonrotating quasi-2D scalar BEC and 2D electrostatics within the GPE and PVM is well established [9]. In this section, we will derive the duality between vortices in a quasi-2D BEC and 2D electrodynamics in general. Motivated by Ref. [39], let us define the vortex charge density $\rho_{v}\left(\bm{r},t\right)$ such that
@@
We define $\rho_{v}\left(\bm{r},t\right)$ in this way because the curl of $\bm{v}_{s}\left(\bm{r},t\right)$ need not equal the Dirac delta function if the superfluid density $n\left(\bm{r},t\right)$ is zero within some finite region around the core of the vortex. For example, the nonlinearity of the field equation for superfluid ${}^{4}\textrm{He}$ changes the curl of the superfluid velocity $\bm{v}_{s}\left(\bm{r},t\right)$ to the smeared delta function (see Eq. (1.64) in Part II of Ref. [34]). Note that setting $\rho_{v}\left(\bm{r},t\right)=\sum_{j=1}^{\infty}q_{j}\left(t\right)\delta\left(\bm{r}-\bm{r}_{\alpha_{j}}\left(t\right)\right)$ corresponds to the PVM, so we also consider beyond the PVM.
@@
However, Eq. (42) does not imply vortex charge conservation; the vortex charge may not be conserved as $q\left(t\right)$ may depend on time $t$. This consideration, including the non-conservation of the vortex charge, is one of the main differences between our work and the previous works relying on the conservation of the vortex charge [32, 69, 76, 77]. In Sec. VI, Eq. (42) will be used to show how to calculate the change of the circulation of $\bm{v}_{s}\left(\bm{r},t\right)$ in time.
@@
If $U_{\rm sf}\left(\bm{r},t\right)$ is constant in space, Eq. (43) cannot be distinguished from Maxwell’s equations with the effective magnetic field strength $\bm{H}_{\rm sf}\left(\bm{r},t\right)=0$. Therefore let us define $\bm{H}_{\rm sf}\left(\bm{r},t\right)$ as
@@
From the Landau’s criterion of superfluidity [78, 57, 11], we will set $c_{\rm sf}$ to be the maximum value of the speed of sound in the scalar BEC, i.e., $c_{s}\coloneqq\sqrt{gn_{\rm max}/M}$, where $n_{\rm max}$ is the maximum value of $n\left(\bm{r},t\right)$.
@@
where the effective free magnetic charge density $\rho_{m,\rm{sf}}\left(\bm{r},t\right)$ is zero [see Eq. (47)]. Equation (51) implies the conservation of the effective free magnetic charge, so the effective free magnetic charge always remains zero.
@@
implying that the vortex (free electric charge) moves parallel to $\bm{v}_{P}\left(\bm{r},t\right)$, not parallel to the superfluid velocity $\bm{v}_{s}\left(\bm{r},t\right)$ in general when the number of vortices is not conserved [55, 56]. The obtained duality is summarized in Fig. 1 and Table 3.
Table 3: Effective Maxwell’s equations for the vortices in the nonrotating quasi-2D scalar BEC. $\rho_{v}\left(\bm{r},t\right)$ is the vortex charge density defined in Eq. (35), $\bm{J}_{\rm sf}\left(\bm{r},t\right)$ is the effective free electric current density defined in Eq. (41), $\bm{J}_{m,\rm{sf}}\left(\bm{r},t\right)$ is the effective free magnetic current density defined in Eq. (49), and $c_{\rm sf}$ is the maximum value of the sound of speed in the scalar BEC. Refer to Eqs. (36), (38), and (44) for the definitions of the effective fields.
@@
We emphasize that our results are of broader generality than those reported in Refs. [38, 39], that assumed a uniform condensate density with negligible fluctuation in a nonrotating system, or Ref. [40], that assumed inhomogeneous time-independent condensate density in a nonrotating system to derive Maxwell’s equations for the (2+1) dimensional superfluid universe. By contrast to these preceding works, the duality described in this Sec. IV can also be applied in the case of an inhomogeneous time-dependent condensate density $n\left(\bm{r},t\right)$ in the (2+1) dimensional spacetime, even in the rotating frame [refer to Eqs. (26) and (28) for a rotating quasi-2D scalar BEC without dissipation]. This extension to the inhomogeneous time-dependent $n\left(\bm{r},t\right)$ is important to study dynamics of vortices because (i) vortices can move, and (ii) $n\left(\bm{r},t\right)=0$ at the core of the vortex. Due to those properties, fluctuation of $n\left(\bm{r},t\right)$ cannot be neglected in general.
@@
The results in this section show that one may consider vortices in a quasi-2D scalar BEC as free electric charges in 2D matter. However, there is one thing missing in the duality we presented: what is the effective magnetic field $\bm{B}_{\rm sf}\left(\bm{r},t\right)$? In the next section, we will define the direction of $\bm{B}_{\rm sf}\left(\bm{r},t\right)$ from the effective Lorentz force and show that the damped PVM [55, 56] can be derived using that effective Lorentz force.
@@ V The effective force acting on vortices
From the duality we found above, one can see that the effective Lorentz force per unit area acting on a vortex, that is, the “effective free electric charge,” has two contributions $\bm{f}_{v}\left(\bm{r},t\right)=\bm{f}_{1}\left(\bm{r},t\right)+\bm{f}_{2}\left(\bm{r},t\right)$, where
@@
Note that Eq. (57) is consistent with the “force per unit length on a vortex line” in Refs. [66, 32]. Another notable thing is that Eq. (60) is similar to the damped-PVM [55, 56]. For a nonrotating quasi-2D scalar BEC confined in a boxlike trapping potential ($V\left(\bm{r},t\right)=0$) as considered in Ref. [56], $M\partial\bm{v}_{s}\left(\bm{r},t\right)/\partial t=\bm{f}_{K}\left(\bm{r},t\right)+\bm{F}_{\rm sf}\left(\bm{r},t\right)-g\nabla n\left(\bm{r},t\right)$, where $\bm{F}_{\rm sf}\left(\bm{r},t\right)$ is the additional force arising from going beyond the nonrotating dissipationless case (refer to Sec. II.2). In the GPE for a nonrotating quasi-2D BEC, minimization of the mean-field energy with respect to the phase of the mean-field $\psi\left(\bm{r},t\right)$ gives $\nabla\cdot\bm{v}_{s}\left(\bm{r},t\right)=0$ [34]. As $\nabla\cdot\bm{P}_{\rm sf}\left(\bm{r},t\right)=0$ in the PVM (see Eq. (40)), for small dissipation where $\nabla\cdot\bm{v}_{s}\left(\bm{r},t\right)\simeq 0$ is still a good approximation, one may set $\bm{P}_{\rm sf}\left(\bm{r},t\right)\simeq c_{1}\left(t\right)\bm{v}_{s}\left(\bm{r},t\right)$, with $c_{1}\left(t\right)$ being some function that only depends on time $t$. Therefore, with a suitable choice of $M_{\rm sf}\left(\bm{r},t\right)$ and an appropriate definition of the effective vortex mass, one may derive the damped-PVM [55, 56]. Or one may use $\bm{P}_{\rm sf}\left(\bm{r},t\right)\simeq c_{1}\left(t\right)\bm{v}_{s}\left(\bm{r},t\right)$ together with the definition of $\bm{P}_{\rm sf}\left(\bm{r},t\right)$ in the second line in Eqs. (38) to get
@@ VI Temporal change of the circulation in a static area
As an example, let us consider the case where only one vortex is positioned at the center of the circular quasi-2D scalar BEC with circular symmetry. We thus assume no external drag or other effect that breaks circular symmetry. Due to this symmetry, $\bm{f}_{K}\left(\bm{r},t\right)$ is perpendicular to the boundary of any disk $\partial\mathcal{D}^{2}_{R}\left(0\right)$ with $R>0$, where $\mathcal{D}^{d}_{R}\left(\bm{r}_{c}\right)$ denotes a $d$-dimensional disk with radius $R$ centered at $\bm{r}_{c}$. Within the GPE, $\bm{F}_{\rm sf}\left(\bm{r},t\right)=0$ for a nonrotating system, while $\bm{F}_{\rm sf}\left(\bm{r},t\right)=M\nabla\left\{\bm{v}_{s}\left(\bm{r},t\right)\cdot\left[\bm{\Omega}_{\perp}\left(\bm{r},t\right)\times\bm{r}\right]\right\}$ in the case of a rotating system (see Eq. (28)), which will also be perpendicular to $\partial\mathcal{D}^{2}_{R}\left(0\right)$ due to the symmetry. Therefore, the vortex will be dynamically stable within the GPE description both in the rotating and nonrotating cases. Conversely, for a rotating or nonrotating quasi-2D scalar circular dissipationless BEC, if the system has circular symmetry, a single vortex cannot emerge at the center if there are no vortices initially. This is consistent with the findings in Ref. [27], where vortex creation is described at the border of the trap.
Note that the results in this section only tells how the circulation (total topological charges of vortices) in a static area changes in time. Therefore, they do not exclude the possibility of the annihilation and creation of two vortices with opposite charges. In the next section, we will show that the stability of the vortex could be understood with the help of the effective Poynting vector introduced in Eq. (52).
VII Stability of vortices in quasi-2D scalar BEC
It is known that the hydrogen atom cannot be stable in classical mechanics due to the radiation. We will show that the vortex number conservation in the PVM for a nonrotating dissipationless quasi-2D scalar BEC in a box trap [42, 43, 44, 45] can be understood by using the duality we constructed in Sec. IV.
In the PVM, the vortex core size is neglected and $\bm{v}_{P}\left(\bm{r},t\right)=\bm{v}_{s}\left(\bm{r},t\right)$, meaning that $\bm{E}_{\rm sf}\left(\bm{r},t\right)=\bm{D}_{\rm sf}\left(\bm{r},t\right)$. Then, regardless of its charge, the vortex velocity is always perpendicular to the effective electric field $\bm{E}_{\rm sf}\left(\bm{r},t\right)$ (see Eqs. (36), (38), and (39)) due to other vortices. Then one can conclude that the vortices cannot collide in the PVM unless one uses the damped PVM or other models that make $\bm{v}_{P}\left(\bm{r},t\right)\neq\bm{v}_{s}\left(\bm{r},t\right)$. This is consistent with the results in [42, 43, 44, 45].
For a nonrotating quasi-2D scalar BEC in a box trap, $n\left(\bm{r},t\right)$ is zero only at the boundary or at the positions of point vortices, and $n\left(\bm{r},t\right)\simeq c$ otherwise where $c$ is some positive constant. Also, if there is no dissipation, $\bm{v}_{s}\left(\bm{r},t\right)$ should be always parallel to the boundary [see Eq. (16)]. This makes $U_{\rm sf}\left(\bm{r},t\right)\simeq\bar{U}_{\rm sf}\left(t\right)$ around any infinitesimal closed curve around the point vortex as long as no other vortices are infinitesimally close to that vortex. Then, in the PVM, the effective Poynting vector $\bm{S}_{\rm sf}\left(\bm{r},t\right)$ is zero around any infinitesimal closed curve around the vortex since vortices cannot collide, meaning that there is no effective radiation and hence the vortex does not lose its energy. In conclusion, the PVM cannot exhibit vortex annihilation/creation in a nonrotating dissipationless quasi-2D scalar BEC in a box trap. Of course, this conclusion does not hold for the damped PVM or other models where $\bm{v}_{P}\left(\bm{r},t\right)\neq\bm{v}_{s}\left(\bm{r},t\right)$, and indeed the damped PVM can explain the vortex annihilation [56]. Also, if there is dissipation, vortices may disappear at the boundary in the PVM since $\bm{v}_{s}\left(\bm{r},t\right)$ is not parallel to the boundary.
From the above discussion, one can infer that the effective photons would be emitted in the annihilation of vortices. As it is known that the phonon emission plays a role in the annihilation and creation of vortices [84, 85], the phonons in a quasi-2D scalar BEC would behave like the effective photons.